Seaborn Co. has identified an investment project with the following cash flows.

Year Cash Flow
1 $800           
2 1,080           
3 1,290           
4 1,120           
Required:
(a)

If the discount rate is 11 percent, what is the present value of these cash flows?

A. 3,205.23 B. 4,290.00 C. 3,354.01 D. 3,638.90 E. 3,278.29
(b)

What is the present value at 18 percent?

A. 4,290.00 B. 3,323.38 C. 2,816.42 D. 2,758.94 E. 3,635.58
(c) What is the present value at 30 percent?
A. 2,194.15 B. 2,903.87 C. 4,290.00 D. 2,233.75 E. 2,274.52
Answer #1

(a)  E. 3,278.29

PV = FV/(1 + r)t

PV@11% = $800/1.11 + $1,080/1.112 + $1,290/1.113 + $1,120/1.114 = $3,278.29

(b) C. 2,816.42

PV = FV/(1 + r)t

PV@18% = $800/1.18 + $1,080/1.182 + $1,290/1.183 + $1,120/1.184 = $2,816.42

(c) D. 2,233.75

PV = FV/(1 + r)t

PV@30% = $800/1.30 + $1,080/1.302 + $1,290/1.303 + $1,120/1.304 = $2,233.75